OpenStudy (anonymous):

Parking in a student lot costs $3 for the first half hour and $1.75 for each hour thereatfer.A partial hour is charged the same as full hour.What is the longest time that a student can park in this lot for $10.00?

OpenStudy (nurali):

cost = 1st hr cost + 1.75*(# of additional hours) C = 3 + 1.75x 10 = 3 + 1.75x 1.75x = 7 x = 4 hrs.